Barn Painting. Melanie can paint a certain barn by herself in x days. Her helper Melissa can paint the same barn by herself in 2x days. Write a rational expression for the fraction of the barn that they complete for 1 day by working together. Evaluate the expression for x = 5.
Answer by edjones(7569)

1/x + 1/2x
=(2+1)/2x LCD=2x
=3/2x of the barn painted in 1 day working together.
.
x=5
3/10 of the barn painted in 1 day working together.
.
Ed